OpenStudy (anonymous):

How many milliliters of a 3.0 M HCl solution are required to make 250.0 milliliters of 1.2 M HCl?

OpenStudy (anonymous):

To solve this, you need to know how many moles of HCl are need. Apply the formula \(n=VM\) to find the moles of HCl needed. \(n\) is moles (mol), \(V\) is the volume (L), and \(M\) is the molarity (M). \(\Rightarrow n_{HCl}=.25 \times 1.2=.3~(mol)\) Next, apply the same formula to find the volume of 3.0M HCl that we need. \(n=VM\Rightarrow V= \Large{\frac{n}{M}} \normalsize =\Large\frac{.3}{3.0} \normalsize =0.1~(L) =100~(mL)\) Or you can use this way to solve the problem. As you know, the moles of HCl is the same. Therefore, \(V_1M_1=V_2M_2 \Longleftrightarrow V_1\times 3.0=250\times 1.2\Longleftrightarrow V_1=100~mL\)